OpenStudy (anonymous):

Will give medals :) Find the inverse of the function. Determine whether the inverse is a function. f(x) = (2x - 3)^2

OpenStudy (anonymous):

solve \[x=(y-3)^2\] for \(y\) in two steps

OpenStudy (anonymous):

but be careful because the first step is \[\pm\sqrt{x}=y-3\] and the \(\pm\) tells you it is not a function

OpenStudy (anonymous):

I honestly don't know how to start the next step.

OpenStudy (anonymous):

well i read the problem incorrectly

OpenStudy (anonymous):

\[x=(2y-3)^2\] so \[\pm\sqrt{x}=2y-3\]

OpenStudy (anonymous):

then add 3 and divide by 2

OpenStudy (anonymous):

\[\pm\sqrt{x}+3=2y\\ y=\frac{\pm\sqrt{x}+3}{2}\]

OpenStudy (anonymous):

That't not an inverse right?

OpenStudy (anonymous):

thanks for your help btw :)

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

what @satellite73 wrote is an inverse, it is just not a function

OpenStudy (anonymous):

Thank you @misty1212 !

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]